OpenStudy (anonymous):

a science teacher has a supply of 5% hydrochloric acid and a supply of 65% hydrochloric acid (HCI). how much of each solution should the teacher mix together to get 42 mL of 45% HCI for an experiment? A- the teacher will need 14 mL of the 5% solution and 28 mL of the 65% solution. B- the teacher will need 28 mL of the 5% solution and 14 mL of the 65% solution. C- the teacher will need 14 mL of the 5% solution and 14 mL of the 65% solution. D- the teacher will need 28 mL of the 5% solution and 28 mL of the 65% solution.

OpenStudy (anonymous):

what do you mean your geting an error

OpenStudy (anonymous):

@Euler271 the other equation which is x+y=42 so solving x+y=42 and 0.05x + 0.65y = 0.45*42 will give u the reqd unique answer

OpenStudy (anonymous):

check if it is A- the teacher will need 14 mL of the 5% solution and 28 mL of the 65% solution.

OpenStudy (anonymous):

matricked is right

OpenStudy (anonymous):

sorry if it took me along time to reply

OpenStudy (anonymous):

i'm gooing to be gone for tell 10:00 or later so sorry about that and i'll get back to you guys and girls

OpenStudy (anonymous):

okay i'm back

OpenStudy (dumbsearch2):

teacher has 2 solutions. 1 = 5% hydrochloric acid. 1 = 65% hydrochloric acid. how much of each solution to get 42 mL of 45% hydrochloric acid? let x = number of mL of first solution. let y = number of mL of second solution. amount of HCI in first solution is .05*x amount of HCI in second solution is .65*y teacher wants 42 mL of 45% HCI. amount of HCI in the solution that the teacher wants is .45 * 42 mL = 18.9 mL. first equation is .05*x + .65*y = 18.9 mL of HCI. first equation tells teacher how much HCI in each solution is required to total the amount of HCI required in the final solution. second equation is x + y = 42 mL of total solution. second equation tells teach how much total of each solution is required to get the proper mix. these 2 equations need to be solved simultaneously to get the right answer. the 2 equations are: .05*x + .65*y = 18.9 x + y = 42 solve for y in the second equation to get y = 42 - x. substitute for y in the first equation to get: .05*x + .65*(42-x) = 18.9 first equation now has 1 unknown and can be solved. solve for x. simplify by removing parentheses in second equation to get: .05*x + .65*42 - .65*x = 18.9 simplify further to get: .05*x + 27.3 - .65*X = 18.9 subtract 27.3 from both sides of the equation to get: .05*x - .65*x = - 8.4 combine like terms to get: -.6*x = -8.4 divide both sides of the equation to get: x = -8.4/-.6 = 14. since x + y = 42, then 14 + y = 42. subtract 14 from both sides of this equation to get: y = 42 - 14 = 28 teacher has x = 14 and y = 28 which means that: teacher needs 14 mL of the 5% solution and 28 mL of the 65% solution. if teacher does that, then teacher gets: .05*14 + .65*28 = .7 + 18.2 = 18.9 mL of HCI. Answer looks good. Answer is: teacher needs 14 mL of 5% solution and 28 mL of 65% solution to get 42mL of 45% solution. 14 + 28 = 42 mL total solution. .05*14 + .65*28 = 18.9 mL of HCI. 18.9 / 42 = .45 = 45% of HCI in final solution.

OpenStudy (dumbsearch2):

So the answer is A. :) Tell me if this helped, and click "Best response" if it did :)

OpenStudy (anonymous):

thank you this helped me alot

OpenStudy (dumbsearch2):

Click "best response" near my answer then :) Thanks.

OpenStudy (anonymous):

there you go man

OpenStudy (anonymous):

well see you in the morning

OpenStudy (dumbsearch2):

Thanks :)

OpenStudy (anonymous):

your welcome buddy

OpenStudy (anonymous):

you now i haven't had this much help since a long time

OpenStudy (anonymous):

bye-bye